### v1

# Learning Concept-Driven Logical Rules for Interpretable and Generalizable Medical Image Classification

Abstract
The pursuit of decision safety in clinical applications highlights the potential of concept-based methods in medical imaging. While these models offer active interpretability, they often suffer from concept leakages, where unintended information within soft concept representations undermines both interpretability and generalizability. Moreover, most concept-based models focus solely on local explanations (instance-level), neglecting the global decision logic (dataset-level). To address these limitations, we propose Concept Rule Learner (CRL), a novel framework to learn Boolean logical rules from binarized visual concepts. CRL employs logical layers to capture concept correlations and extract clinically meaningful rules, thereby providing both local and global interpretability. Experiments on two medical image classification tasks show that CRL achieves competitive performance with existing methods while significantly improving generalizability to out-of-distribution data. The code of our work is available at https://github.com/obiyoag/crl.
Keywords:
Explainable-AI Concept Learning Rule-based Model.
1 Introduction
Deep learning models, especially those operating as black boxes, have shown great promise in medical imaging applications [10, 13, 12]. However, the high standards of trust and accountability required in healthcare have spurred growing interest in explainable approaches [22]. Merely attempting to interpret black-box models could inadvertently reinforce flawed clinical practices and lead to harmful medical outcomes [19]. Therefore, it is essential to develop self-explanatory methods for medical imaging. Recently, Concept Bottleneck Models (CBMs) [11] have gained attention as an effective approach to foster active interpretability [25]. In CBMs, a set of interpretable intermediate concepts is first generated by a concept predictor, and these concepts are then used to make the final predictions by a label predictor. Building on CBMs, Concept Embedding Models (CEMs) [6] improve the predictive performance by employing high-dimensional concept representations and demonstrate their applicability in medical domain [7].
However, both CBMs and CEMs suffer from concept leakage [14], where the label predictor inadvertently exploits unintended information from concept probabilities or embeddings. This leakage compromises both interpretability and generalizability [15]. In terms of interpretability, the predictions are influenced not only by the intended concepts but also by image information encoded in the soft concepts (i.e. probabilities or embeddings). The concept predictor no longer needs to faithfully predict the concepts to obtain accurate label predictions. Regarding generalizability, reliance on unintended leaked information may lead to overfitting, rendering the models less robust to distribution shifts. The overfitting occurs because the soft concepts may encode spurious correlations that are not representative of the true underlying concepts.
Furthermore, most concept-based models only provide explicit local explanations by focusing on individual predictions, detailing how predicted concepts influence each decision [16, 24]. They offer insights for particular instances but may not fully capture the overall decision logic across the entire category or dataset. To achieve a more comprehensive interpretability, it’s essential to integrate both local and global explanations [25, 20]. A recent work, named Deep Concept Reasoner (DCR) [2], explores to extract syntactic logical rules from concept embeddings. However, building upon fuzzy logic and concept embeddings, DCR’s decision-making process becomes less transparent and susceptible to concept leakage.
To mitigate the issue of concept leakage and unify local and global explanations, we propose the Concept Rule Learner (CRL), a framework to learn Boolean logical rules from visual concepts. To alleviate concept leakage, CRL employs binary concept values (i.e., 0 and 1), which inherently carry less unintended image information. CRL introduces logical layers [23] to model Boolean logical operations, thereby capturing correlations between concepts. Each logical layer comprises a conjunction layer and a disjunction layer, which perform the AND and OR operations, respectively. The connections across these layers generate flexible decision rules, and the final prediction for an input image is derived by aggregating the contributions of triggered rules through linear weights. By making decisions based on domain-invariant logical rules, CRL not only incorporates global interpretability, but also improves generalizability to unseen domains. Our main contributions can be summarized as follows:
  • $\bullet$
    We propose CRL, a novel framework that learns Boolean logical rules from binarized visual concepts to model concept correlations, while mitigating the issue of concept leakage.
  • $\bullet$
    CRL not only delivers concept explanations for individual predictions but also extracts decision rules for the entire datasets, thereby unifying local and global interpretations.
  • $\bullet$
    We evaluate the effectiveness of the proposed methods on two medical image classification tasks. The experimental results demonstrate that our approach can extract meaningful concept logical rules and exhibits superior generalizability to the unseen dataset.
2 Method
2.1 Model Architecture
Figure 1: The architecture of Concept Rule Learner (CRL). CRL is a composition of four functions $(f\circ r\circ q\circ g)$, where $g(\cdot)$ represents the concept predictor, $r(\cdot)$ denotes a stack of logical layers, $q(\cdot)$ means discretization and $f(\cdot)$ corresponds to a linear layer. The decision interpretation is presented in the dashed box, illustrating how concepts and rules contribute to the final prediction.
Given an image $\mathbf{x}\in\mathbb{R}^{n}$ annotated with a task label $y\in\{1,\cdots,L\}$ and $K$ concept labels $\mathbf{c}\in\{0,1\}^{K}$, our objective is to develop a Concept Rule Learner (CRL) capable of performing medical image classification through concept-driven logical rules. As illustrated in Fig. 1, the CRL framework is a quadruple of functions $\boldsymbol{(g,q,r,f)}$, whose composition $(f\circ r\circ q\circ g)$ predicts the label based on the derived concepts and rules. The first function $\boldsymbol{g}:\mathbb{R}^{n}\to[0,1]^{K}$, referred to as the concept predictor, learns a mapping from the input image $\mathbf{x}$ to a set of concept activations $\hat{\mathbf{c}}=g(\mathbf{x})\in[0,1]^{K}$, where $\hat{c}_{i}$ represents the probability supporting the presence of the $i$-th concept. The second function $\boldsymbol{q}:[0,1]^{K}\to\{0,1\}^{K}$ serves as a binarizer that converts the concept activations into binary values. The third function $\boldsymbol{r}:\{0,1\}^{K}\to\{0,1\}^{R}$, consists of a series of logical layers adopted from [23]. This function maps the activated concepts to a set of rule activations $\mathbf{r}=r(q(\hat{\mathbf{c}}))\in\{0,1\}^{R}$, where $R$ denotes the total number of extracted logical rules. A rule activation $r_{i}=1$ indicates that the $i$-th rule is activated for the input image $\mathbf{x}$ and 00 for otherwise. The fourth function $\boldsymbol{f}:\{0,1\}^{R}\to\mathbb{R}^{L}$, represents the final linear layer, which learns the mapping from the activated logical rules to the task logits $\hat{\mathbf{y}}=f(\hat{\mathbf{r}})\in\mathbb{R}^{L}$.
Figure 2: (a) An example of two adjacent logical layers, where the directed arrows indicates the presence of the connections between nodes. The corresponding rules could be derived by analyzing the connections. (b) The computation graph of CRL. Arrows with solid lines represent forward pass while arrows with dashed lines represent backpropagation.
2.2 Logical Operation Modelling
To learn logical rules from concept activations, the function $r(\cdot)$ adopts a series of logical layers to model the concept-based rules with Boolean logic. Let $\mathcal{N}^{l}$ denote the $l$-th logical layer, where $n_{j}^{l}$ represents the $j$-th node within the layer. The output of layer $\mathcal{N}^{l}$ is a vector containing the values of all the nodes, denoted as $\mathbf{n}^{l}$. As shown in Fig. 2(a), each logical layer $\mathcal{N}^{l}$ consists of a conjunction layer and a disjunction layer, denoted by $\mathcal{A}^{l}$ and $\mathcal{B}^{l}$, respectively. The $i$-th node in the conjunction layer, denoted as $a_{i}^{l}$, represents the conjunction (AND operation) of nodes from the preceding layer connected to it. Conversely, the $i$-th node in the disjunction layer, denoted as $b_{i}^{l}$, encapsulates the disjunction (OR operation) of its connected predecessors. Formally, the nodes in the conjunction and disjunction layers are defined as:
$a_{i}^{(l)}=\bigwedge_{W_{i,j}^{(l,0)}=1}n_{j}^{l-1},\;\;\;\;\;\;b_{i}^{l}=\bigvee_{W_{i,j}^{(l,1)}=1}n_{j}^{l-1},$ (1)
where $W^{(l,0)}$ denotes the adjacency matrix of the conjunction layer $\mathcal{A}^{l}$ and the previous layer $\mathcal{N}^{l-1}$, with $W_{i,j}^{(l,0)}\in\{0,1\}$. Specifically, $W_{i,j}^{(l,0)}=1$ indicates presence of an edge connecting $a_{i}^{l}$ to $n_{j}^{l-1}$, while $W_{i,j}^{(l,0)}=0$ indicates the absence. Similarly, $W^{(l,1)}$ is the adjacency matrix of the disjunction layer $\mathcal{B}^{l}$ and $\mathcal{N}^{l-1}$. These adjacency matrices are treated as the weight matrices for the logical layers, analogous to weight matrices in neural networks. The output of the $l$-th layer is the concatenation of the outputs of the conjunction and disjunction layers, i.e., $\mathbf{n}^{l}=\mathbf{a}^{l}\oplus\mathbf{b}^{l}$, where $\mathbf{a}^{l}$ and $\mathbf{b}^{l}$ are the outputs of $\mathcal{A}^{l}$ and $\mathcal{B}^{l}$ respectively. Fig. 2(a) illustrates an example of two adjacent logical layers, where the directed arrows indicates the presence of the connections between nodes. By examining the weights $W^{(l,0)}$ and $W^{(l,1)}$, we can derive the corresponding rules in both conjunctive and disjunctive normal forms.
2.3 Training Paradigm
While the logical layers are capable of expressing Boolean operations, their non-differentiable structure makes CRL challenging to optimize. To address this issue, we introduce continuous logical layers in training. These layers are differentiable and share the same parameters as the discrete counterparts, enabling end-to-end optimization while preserving the interpretability of the original design.
Let $\bar{W}^{(l,0)},\bar{W}^{(l,1)}\in[0,1]$ denote the continuous weight matrices of conjunction and disjunction layers. To make Eq. 1 differentiable, we leverage the logical activation functions introduced by [18]:
$\textit{Conj}(\mathbf{n},\bar{W}_{i})=\mathbb{P}(\prod_{j=1}^{N}F_{c}(n_{j},\bar{W}_{i,j})),\;\;\textit{Disj}(\mathbf{n},\bar{W}_{i})=1-\mathbb{P}(\prod_{j=1}^{N}F_{d}(n_{j},\bar{W}_{i,j})),$
where $N$ is the number of nodes, $F_{c}(n,w)=1-w(1-n)$ and $F_{d}(n,w)=1-n\cdot w$. If $\mathbf{n}$ and $W_{i}$ are both binary vectors, then $\textit{Conj}(\mathbf{n},W_{i})=\bigwedge_{W_{i,j}=1}\mathbf{n}_{j}$ and $\textit{Disj}(\mathbf{n},W_{i})=\bigvee_{W_{i,j}=1}\mathbf{n}_{j}$. $\mathbb{P}(x)=1/(1-\log x)$ is a projection function to prevent gradients vanishing. After using continuous weights and logical activation functions, the nodes in continuous logical layers are defined as follows:
$\bar{a}_{i}^{l}=\textit{Conj}(\bar{\mathbf{n}}^{l-1},\bar{W}_{i}^{(l,0)}),\;\;\;\;\;\;\bar{b}_{i}^{l}=\textit{Disj}(\bar{\mathbf{n}}^{l-1},\bar{W}_{i}^{(l,1)}).$
By employing continuous logical layers, CRL facilitates end-to-end training. The computational graph of CRL is illustrated in Fig. 2(b). As shown, the concept predictor generates concept activations $\hat{\mathbf{c}}=g(\mathbf{x})$ from input images. The binarized concept activatioes are passed to both the discrete and continuous logical layers to generate the task predictions given by $\hat{\mathbf{y}}=f(r(q(\hat{\mathbf{c}}))|W)$ and $\bar{\mathbf{y}}=f(r(q(\hat{\mathbf{c}}))|\bar{W})$. The objective function is defined as:
$\mathcal{L}=\mathcal{L}_{t}(\hat{\mathbf{y}},y)+\mathcal{L}_{c}(\hat{\mathbf{c}},\mathbf{c})+\lambda||\bar{W}||_{2},$ (2)
where $\mathcal{L}_{t}$ is the cross-entropy loss, $\mathcal{L}_{c}$ represents the mean cross-entropy loss across all training concepts and $\lambda$ is the regularization hyperparameter that controls the complexity of the logical layers. For backpropagation, as indicated by the dotted arrow in Fig. 2(b), the gradients $\partial\mathcal{L}_{t}/\partial\hat{\mathbf{y}}$ are grafted onto the backward pass of $f(r(\cdot|\bar{W}))$. To address the non-differentiability of the binarizer, we utilize a Straight-Through Estimator (STE) [9], which approximates the gradients for the discretization.
2.4 Decision Interpretation
As illustrated in the dashed box of Fig. 1, CRL provides concept explanations for individual predictions as well as rule explanations for the entire dataset, thereby integrating both local and global explanations. After training, the weights of logical layers $r(\cdot)$ are analyzed to summarize $R$ logical rules, where $R$ depends on the number of nodes and the hyperparameter $\lambda$ in Eq. 2. During inference, the concept predictor $g(\cdot)$ first extracts concepts from images and matches them against the learned rules. Each rule is associated with a set of class-specific weights from the linear layer $f(\cdot)$. When a rule is triggered by a match (bold in Fig. 1), its corresponding weights are added to the overall class logits $\hat{\mathbf{y}}$. The final prediction is determined by the cumulative logits, which aggregate the contribution of all matched rules.
3 Experiments
3.1 Experimental Setup
3.1.1 Tasks and datasets:
To evaluate the proposed method across various medical scenarios, we assess two tasks: skin disease diagnosis and white blood cell (WBC) classification. Skin disease diagnosis: We employ the Fitzpatrick17k (F17k) dataset [8] and the Diverse Dermatology Images (DDI) dataset [4], incorporating concept annotations from the SkinCon dataset [5]. The SkinCon dataset comprises 48 concepts annotated by board-certified dermatologists. Following the approach in [17], we restrict our training and testing to images labeled as Benign (Ben) or Malignant (Mal). WBC classification: We utilize the PBC dataset [1], along with the concept annotations from the WBCAtt dataset [21]. The WBCAtt dataset contains 24 morphological attributes, and the classification includes five distinct classes.
3.1.2 Implementation details:
We employ ResNet-34 pretrained on the ImageNet dataset, as the backbone network. The balance parameter between concept and diagnosis loss is set to 1. We use the AdamW optimizer with an initial learning rate of $5\times 10^{-5}$ and a weight decay of 0.01. The learning rate decays to zero following a cosine scheduler. Models are trained for 300 epochs with a batch size of 64. CRL is implemented with two logical layers, each comprising 256 nodes. The hyperparameter $\lambda$ set to $5\times 10^{-6}$ to control the complexity of the rules. For evaluation, we report the average accuracy (ACC) and F1 score (F1) for both concept prediction and diagnosis tasks. For skin disease diagnosis, we perform 5-fold cross-validation. For WBC classification, we adopt the original data split from [21], conducting experiments with three different random seeds.
3.2 Results
Table 1: Performance comparison on skin disease diagnosis and WBC classification tasks. Bold text indicates the best results, while underlined text denotes the second-best results. The symbol $\star$ denotes methods employ binary concept values, while $\dagger$ indicates methods with global interpretability.
Dataset Method Concept Metric Diagnosis Metric
ACC(%) F1(%) ACC(%) F1(%)
F17k CBM [11] 91.41±0.19 59.44±0.97 76.37±3.45 76.30±3.45
align-CBM [17] 89.77±0.52 58.98±1.41 75.93±2.39 75.83±2.41
CEM [6] 91.85±0.55 59.11±2.06 76.26±2.59 76.17±2.52
evi-CEM [7] 92.04±0.80 58.65±1.71 76.47±1.69 76.39±1.66
hard-CBM 86.36±1.09 49.57±0.48 73.51±2.94 73.43±2.92
DCR [2] 91.03±0.80 49.40±1.00 75.05±2.12 74.97±2.12
[…]
PBC CBM [11] 95.21±0.10 91.65±0.15 98.93±0.14 98.44±0.22
align-CBM [17] 95.01±0.23 89.95±0.59 99.14±0.05 99.25±0.52
CEM [7] 94.98±0.34 90.86±0.69 99.43±0.12 99.23±0.18
evi-CEM [7] 94.44±0.91 89.18±2.09 99.57±0.03 99.42±0.03
hard-CBM 64.53±2.42 56.24±2.74 98.22±0.48 97.52±0.73
DCR [2] 92.79±0.43 82.92±0.64 98.93±0.32 98.47±0.49
[…]
3.2.1 Model utility analysis:
To showcase the classification utility of the proposed method, we compare CRL with other concept-based methods on F17k and PBC datasets. The comparative methods include CBM [11], align-CBM [17], CEM [6], evi-CEM [7] and DCR [2]. Among these, align-CBM integrates clinical knowledge to prioritize the most relevant, while evi-CEM employs evidential learning to model concept uncertainty. Hard-CBM, which serves as a baseline, only accepts binary concept values without logical layers. The comparison results are presented in Table 1. From the results, we observe that CRL achieves comparable predictive performance with other CBM variants, though it exhibits a performance trade-off compared to methods based on concept embeddings. Notably, when comparing CRL and hard-CBM, both of which utilize binary concepts, CRL significantly outperforms hard-CBM. This improvement can be attributed to the logical layers in CRL, which capture the concept correlations with logical operations.
3.2.2 Model generalizability analysis:
Table 2: Performance comparison on the unseen DDI dataset. Bold text indicates the best results, while underlined text denotes the second-best results. The symbol $\star$ denotes methods employ binary concept values, while $\dagger$ indicates methods with global interpretability.
[…]
Concept Metric Diagnosis Metric Concept Metric Diagnosis Metric
ACC(%) F1(%) ACC(%) F1(%) ACC(%) F1(%) ACC(%) F1(%)
CBM [11] 90.59±0.27 53.09±0.44 62.36±4.05 57.49±2.02 0.82 6.35 14.01 18.81
align-CBM [17] 89.34±0.25 51.47±0.22 63.13±4.01 59.66±2.18 0.33 7.51 12.80 16.17
CEM [6] 91.66±0.26 53.21±0.57 62.39±2.59 57.42±2.01 0.19 5.90 13.87 18.75
evi-CEM [7] 91.35±0.20 52.17±0.64 64.25±2.04 58.82±1.62 0.69 6.48 12.12 17.48
hard-CBM 84.63±1.20 48.48±0.28 63.77±2.35 56.96±1.52 1.73 1.09 9.74 16.47
DCR [2] 90.97±0.49 47.89±0.72 62.23±4.09 57.50±2.35 0.06 1.51 12.82 17.47
[…]
To evaluate generalizability, we assess the out-of-domain (OOD) performance of the models on the unseen DDI dataset, using models trained on the source F17k dataset. As reported in Table 2, we can observe that concept models relying on soft concepts (probabilities or embeddings) exhibit relatively large performance drops, whereas methods employing binary concept values experience smaller declines. This suggests that binary concepts help mitigate concept leakage. Notably, although both CRL and hard-CBM utilize binary concepts, CRL achieves a diagnostic ACC of 73.46, outperforming the second-best method by 9.21 while exhibiting the smallest performance drop. This improvement can be attributed to the domain-invariant logical rules extracted by logical layers, which enhance robustness to distribution shifts.
3.2.3 Model interpretability analysis:
Figure 3: An illustration of concept logical rules for both skin disease diagnosis task (left) and WBC classification (right). Note: We only present the rule weights and logits of Neutrophil (Neut) and Lymphocyte (Lymph) for WBC classification.
To illustrate that CRL could generate meaningful logical rules, we present the rules obtained from both skin disease diagnosis and WBC classification tasks, as illustrated in Fig. 3. For skin disease diagnosis, we observe that concepts Telangiectasia, Ulcer and Crust are associated with Malignant, aligning with established clinical knowledge [3]. In case of WBC classification, the rules indicate that Lymphocytes are characterized by a High NC ratio and a Round nucleus, while Neutrophils typically exhibit Pink granules and Dense chromatin, consistent with clinical observations [21]. The case studies demonstrate that CRL can effectively extract concept-based logical rules that are clinically meaningful.
4 Conclusion
This paper proposes CRL, a framework for interpretable medical image classification that mitigates concept leakage and unifies local and global explanations. By leveraging binary concepts and learnable logical layers, CRL effectively models concept correlations and extracts clinically meaningful decision rules. Experiments on two tasks demonstrate that CRL not only achieves competitive performance but also exhibits superior generalizability to unseen data.